# Stabilizing Non-Abelian Topological Order against Heralded Noise via Local Lindbladian Dynamics

### v1

@@ C.3 Stabilizer tableaux
Having identified the form of instantaneous stabilizers, we now present the method used in our simulations for encoding them efficiently. The series of steps discussed in the previous subsection only switch the signs $\mathsf{s_{v}}$ associated with the vertex stabilizers $q_{v}=\mathsf{s_{v}}B_{v}$ without modifying the form of the operator. Hence, the instantaneous state of the vertex stabilizers can be represented by a length $2L^{2}$ array whose elements keep track of these signs.
The remaining plaquette type stabilizers in $\mathcal{G}$ are modified in non-trivial manner which includes multiple $A$ operators with additional Pauli-Z decorations in given stabilizer. More explicitly, a generator $g_{j}\in\mathcal{G}$ can be written in terms of its constituents as
@@
where $N_{q}=3L^{2}$ and $N_{p}=L^{2}$ is the total number of qubits and plaquettes in the system respectively. Here the products run over all three colors $r,g,b$ but we suppress the color index for brevity. The variables $\mathsf{z_{j,k}}$ and $\mathsf{a_{j,k}}$ are either $1$ or 00 depending on whether the associated operator is present or absent respectively and $\mathsf{s_{j}}=\pm 1$ is overall sign. As discussed in Eq. (57), there are $L^{2}-3=N_{p}-3$ independent generators. We collect all of these operator representations in a $(N_{p}-3)\times(N_{q}+N_{p})$ dimensional matrix
@@
where each entry is either 00 or $1$. The overall sign $\mathsf{s_{j}}$ is stored in a separate length $N_{p}-3$ array. The stat of destabilizers can be analogously tracked using matrix $\mathsf{D}$ where $j^{th}$ row represents the instantaneous operator $d_{j}$.
A key feature of this representation is that the matrices $\mathsf{G}$ and $\mathsf{D}$ can be updated efficiently. The update after application of $X$ or $Z$ to a qubit has a non-trivial effect on the stabilizer only if the plaquette sharing that qubit is present in its operator string. Hence, such an update only requires scanning through two columns of the matrix to determine the modified stabilizers. The updates are done by adding new $Z$ decorations or changing sign which involve updating only $O(1)$ entries per row. The implementation of stabilizer measurement requires determination of the commutation relation between $A_{p}$ and a row of $\mathsf{G}$ matrix. The operator $A_{p}$ has non-trivial commutation relation only with 6 nearby plaquettes according to Eq. (22). Hence, the checks similar to Eq. (65) to determine if we are in case (a) or case (b) discussed above require scanning through $6$ columns of matrix per measurement event and has $O(L^{2})$ operation cost. The simulation of deterministic measurement outcome (see case b) has an additional overhead of generating matching graph that pairs all defects. We determine these pairings using pyMatching package that solves this problem in $poly(L)$ time [102]. The part of update rules that depend on algebraic relations between stabilizer operators $A,B$ and Pauli operators are pre-computed at the beginning of the simulation. This includes indices of qubits coupled to given qubit by $CZ$ gates and indices of vertex operators that reside in the interior of overlapping plaquette. Since this is still $O(1)$ data per qubit/plaquette, it does not change overall memory requirement to simulate the given system. The locations of $CZ$ gates in the logical $X$ string are also pre-computed which incurs similar cost ($O(L)$ for each of the three $X$-logicals).
@@ Appendix D Mean field equations for erasure dynamics
Here, the first term corresponds to the addition of $X$ flags at unoccupied sites after an erasure error. The second term models the contribution from the leaf-correction protocol where an $X$ flag on a given edge is removed if it is the only flagged edge incident on one of its vertices. We note that the loop moves in the main text were introduced to remove closed loops of flags that are otherwise unaffected by the leaf correction operations. Since the mean field approximation neglects any inter-site correlations, it does not distinguish between leaf and loop configurations. Thus, the correction term in Eq. (74) is sufficient to remove all flags in this approximation and we neglect the contributions from loop moves to obtain a simplified description.
@@
Figure 20: Connected clusters of flags in active phase: The largest connected cluster of flags is determined in each Monte Carlo trajectory after $500$ sweeps evaluated at representative point in the active phase $\gamma_{x}=35,\gamma_{z}=500$. (top panel) Probability density $\rho(s)$ that the largest cluster in given Monte Carlo realization is of size $s$ is plotted for various system sizes for both $X$ and $Z$ type flags. The size $s$ of a cluster is defined as total number of flags in the connected cluster. The data is shown for $1000$ samples. (bottom panel) The average size of largest cluster $\langle S^{\text{max}}\rangle$ obtained from the histograms is plotted as a function of increasing system size $L$. The error bars represent the standard error in the average value and dashed lines are obtained by fitting the data points to function $a\log{(L/L_{0})}$.

### v2

@@ C.3 Stabilizer tableaux
Having identified the form of instantaneous stabilizers, we now present the method used in our simulations for encoding them efficiently. The series of steps discussed in the previous subsection only switches the signs $\mathsf{s_{v}}$ associated with the vertex stabilizers $q_{v}=\mathsf{s_{v}}B_{v}$ without modifying the form of the operator. Hence, the instantaneous state of the vertex stabilizers can be represented by a length $2L^{2}$ array whose elements keep track of these signs.
The remaining plaquette-type stabilizers in $\mathcal{G}$ are modified in a non-trivial manner, which includes multiple $A$ operators with additional Pauli-Z decorations in the given stabilizer. More explicitly, a generator $g_{j}\in\mathcal{G}$ can be written in terms of its constituents as
@@
where $N_{q}=3L^{2}$ and $N_{p}=L^{2}$ is the total number of qubits and plaquettes in the system, respectively. Here, the products run over all three colors $r,g,b$ but we suppress the color index for brevity. The variables $\mathsf{z_{j,k}}$ and $\mathsf{a_{j,k}}$ are either $1$ or 0 depending on whether the associated operator is present or absent respectively and $\mathsf{s_{j}}=\pm 1$ is the overall sign. As discussed in Eq. (58), there are $L^{2}-3=N_{p}-3$ independent generators. We collect all of these operator representations in a $(N_{p}-3)\times(N_{q}+N_{p})$ dimensional matrix
@@
where each entry is either 0 or $1$. The overall sign $\mathsf{s_{j}}$ is stored in a separate length $N_{p}-3$ array. The state of destabilizers can be analogously tracked using matrix $\mathsf{D}$ where $j^{th}$ row represents the instantaneous operator $d_{j}$.
A key feature of this representation is that the matrices $\mathsf{G}$ and $\mathsf{D}$ can be updated efficiently. The update after application of $X$ or $Z$ to a qubit has a non-trivial effect on the stabilizer only if the plaquette sharing that qubit is present in its operator string. Hence, such an update only requires scanning through two columns of the matrix to determine the modified stabilizers. The updates are done by adding new $Z$ decorations or changing the sign, which involves updating only $O(1)$ entries per row. The implementation of stabilizer measurement requires determination of the commutation relation between $A_{p}$ and a row of $\mathsf{G}$ matrix. The operator $A_{p}$ has non-trivial commutation relations only with 6 nearby plaquettes according to Eq. (23). Hence, the checks similar to Eq. (66) to determine if we are in case (a) or case (b) discussed above require scanning through $6$ columns of the matrix per measurement event and has $O(L^{2})$ operation cost. The simulation of deterministic measurement outcome (see case b) has an additional overhead of generating matching graph that pairs all defects. We determine these pairings using pyMatching package that solves this problem in $poly(L)$ time [106]. The part of the update rules that depend on algebraic relations between stabilizer operators $A,B$ and Pauli operators are pre-computed at the beginning of the simulation. This includes indices of qubits coupled to a given qubit by $CZ$ gates and indices of vertex operators that reside in the interior of an overlapping plaquette. Since this is still $O(1)$ data per qubit/plaquette, it does not change the overall memory requirement to simulate the given system. The locations of $CZ$ gates in the logical $X$ string are also pre-computed, which incurs a similar cost ($O(L)$ for each of the three $X$-logicals).
@@ Appendix D Mean field equations for erasure dynamics
Here, the first term corresponds to the addition of $X$ flags at unoccupied sites after an erasure error. The second term models the contribution from the leaf-correction protocol, where an $X$ flag on a given edge is removed if it is the only flagged edge incident on one of its vertices. We note that the loop moves in the main text were introduced to remove closed loops of flags that are otherwise unaffected by the leaf correction operations. Since the mean field approximation neglects any inter-site correlations, it does not distinguish between leaf and loop configurations. Thus, the correction term in Eq. (75) is sufficient to remove all flags in this approximation, and we neglect the contributions from loop moves to obtain a simplified description.
@@
Figure 20: Connected clusters of flags in active phase: The largest connected cluster of flags is determined in each Monte Carlo trajectory after $500$ sweeps evaluated at a representative point in the active phase $\gamma_{x}=35,\gamma_{z}=500$. (top panel) Probability density $\rho(s)$ that the largest cluster in a given Monte Carlo realization is of size $s$ is plotted for various system sizes for both $X$ and $Z$ type flags. The size $s$ of a cluster is defined as the total number of flags in the connected cluster. The data is shown for $1000$ samples. (bottom panel) The average size of the largest cluster $\langle S^{\text{max}}\rangle$ obtained from the histograms is plotted as a function of increasing system size $L$. The error bars represent the standard error in the average value, and dashed lines are obtained by fitting the data points to function $a\log{(L/L_{0})}$.